OpenStudy (anonymous):

can someone hell me please Choose the correct description of the graph of the compound inequality 3x + 1 > 10 and 4x less than or greater to 20. A number line with an open circle on 3, a closed circle on 5, and shading in between. A number line with an open circle on 3, shading to the left, and a closed circle on 5, shading to the right. A number line with a closed circle on 3, shading to the left, and an open circle on 5, and shading to the right. A number line with a closed circle on 3, an open circle on 5, and shading in between

OpenStudy (anonymous):

3x + 1 > 10 4x <= 20 3x > 10 - 1 x <= 20/4 3x > 9 x <= 5 x > 3 Learn this... if there is an equal sign, it is a closed circle if there is no equal sign, it is an open circle if < less then, shading to the left if > greater then, shading to the right answer is A : open circle on 3,(because there is no equal sign), closed circle on 5,(because there is an equal sign), shading in between.(because x is greater then 3, shading goes to the right, and because x is less then 5, shading goes to the left.

OpenStudy (anonymous):

thank you so much (: I appreciate it